OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

if the point are (a,b) and(c,d) then the equation of line is (y-b)(c-a)=(x-a)(d-b) put the values of a, b ,c. d and get the answer :)

OpenStudy (anonymous):

x+7y=12 is the equation according to me..

OpenStudy (anonymous):

thats what i got also Vjha

OpenStudy (anonymous):

so it means you are right.

OpenStudy (anonymous):

thank u very much